1/2

Jean de La Fontaine,  Jean de La Fontaine,
The Cricket and the Ant  La Cigale et la Fourmi

tradukita de Donald P. Webb

The cricket had sung her song all summer long but found her victuals too few when the north wind blew. Nowhere could she espy a single morsel of worm or fly. Her neighbor, the ant, might, she thought, help her in her plight, and she begged her for a little grain till summer would come back again. “By next August I’ll repay both Interest and principal; animal’s oath.” Now, the ant may have a fault or two But lending is not something she will do. She asked what the cricket did in summer. “By night and day, to any comer I sang whenever I had the chance.” “You sang, did you? That’s nice. Now dance.”

La cigale, ayant chanté Tout l’été, Se trouva fort dépourvue Quand la bise fut venue. Pas un seul petit morceau De mouche ou de vermisseau Elle alla crier famine Chez la fourmi sa voisine, La priant de lui prêter Quelque grain pour subsister Jusqu’à la saison nouvelle �Je vous paierai, lui dit-elle, Avant l’oût, foi d’animal, Intérêt et principal .� La fourmi n’est pas prêteuse ; C’est là son moindre défaut �Que faisiez-vous au temps chaud? Dit-elle à cette emprunteuse. Nuit et jour à tout venant Je chantais, ne vous déplaise.

- Vous chantiez? j’en suis fort aise. Eh bien : dansez maintenant.�

…  …

2/2

Traduko de la Franca poemo “La Cigale et la Fourmi” de Jean de La Fontaine (∗1621-07.08 – †1695-04.12) en la Anglan de Donald P. Webb (∗1939).

Arg-549-1111 (2010-08-06 19:10:14)

Mi, Manfred Retzlaff, transprenis tiun ĉi tradukon el la retejo http: // www. bewilderingstories. com/ issue209/ cigale. html . La tradukinto, s-ro Don Webb, permesis tion mi. - I, Manfred Retzlaff, found this translation in the website http: // www. bewilderingstories. com/ issue209/ cigale. html . The author, Mr. Don Webb, gave me the permission.

Verkinto de tiu ĉi Franca poemo estas Jean de La Fontaine (∗1621-07.08 – †1695-04.12).

Arg-549-1103 (2010-08-07 13:53:59)

Mi, Manfred Retzlaff, kopiis tiun ĉi tekston el http: // www. lafontaine. net/ lesFables/ afficheFable. php? id= 1 .

http://www.poezio.net/version?poem-id=549&version-id=,1111,1103